# ComplexityMeasures.jl: scalable software to unify and accelerate entropy and complexity timeseries analysis

### v1

@@ 2.3 Clarifying and educative approach to naming
The literature is full of complexity measures that are named similarly, but represent fundamentally different concepts. For example, the term "bubble entropy" [18] is not an entropy per se, but a scaled difference between two Rényi [43] entropies which have been computed by discretizing the input data in a particular manner which has to do with the bubble sort algorithm. Rényi entropy, however, is an axiomatically well-founded entropy. Similarly, the permutation entropy [10] or wavelet entropy [11] are not distinct forms of entropy in terms of the mathematical definition of entropy. They are just the Shannon entropy computed by discretizing the input data into outcomes based on permutation patterns or wavelet coefficients, respectively. It is in fact common that papers that introduce new discretization procedures, i.e., outcome spaces, are given the name “entropy” [44, 16, 15].
We believe that this abuse of terminology can be confusing, especially to newcomers to the field. Indeed, while teaching nonlinear timeseries analysis we experience that students often interpret the Shannon entropy and permutation entropy as two different quantities. With ComplexityMeasures.jl and this paper, we aim to clarify this naming confusion, while also highlighting what the common elements between different complexity measures are: estimating probabilities from data. That is why in the software we do not promote function names like “permutation entropy”. That being said, we understand that some terms (like the permutation entropy) are very well recognized in the field and should be accessible. Thus, for a few handpicked complexity measures we provide a shorter syntax, such as the function entropy_permutation(input; m = 3). Nevertheless we make it clear in the documentation of these “convenience functions” that they do not provide a genuinely new entropy even if named as such.
@@ 2.4 Software quality
ComplexityMeasures.jl was implemented following best practices in scientific code [38]. The software is free and open source (MIT-licensed), hosted on GitHub, and also available through the Julia package manager. The repository is composed in total of 3,834 lines of source code, 2,406 lines of test code, and 5,984 lines of documentation text according to PackageAnalyzer.jl [45]. The online documentation is very extensive. It features a central tutorial; full API reference listing outcome spaces, entropy/complexity measures, and estimators provided by the software; more than a dozen of individual examples that are created by showing real code tied with its output; a developer’s documentation for contributing more features to the library. The documentation also cites all research articles introducing the implemented measures/estimators via BiBTeX, and provides an explanation and/or description of each measure/estimator implemented, making it straightforward to understand what this measure is and what it estimates without needing to consult the research article. The software is extensively tested, with coverage of 9̃0%, which means that at least 90% of the source code lines are explicitly called in the test suite. Both tests and documentation are run through continuous integration upon every committed change to the software. ComplexityMeasures.jl is continuously improved and follows agile development practices [46]. A new feature, or even the smallest bugfix, is immediately released as a new software version, which the users can obtain instantly via a standard update command provided by the Julia language. The quoted numbers in this subsection refer to version v3.5 of the software.
@@ 2.5 Part of a greater whole
ComplexityMeasures.jl can be used as a standalone software. However, it is also part of a greater whole. It is a component of the DynamicalSystems.jl [47] software library for nonlinear dynamics and nonlinear timeseries analysis, and is also the basis for the CausalityTools.jl library for relational (or causal) timeseries analysis [39]. In this way, ComplexityMeasures.jl integrates with a wider ecosystem for timeseries and data analysis.
@@
For example, it can be immediately used with TimeseriesSurrogates.jl [48] to perform surrogate analysis testing for nonlinearities, thus eliminates the need to re-implement surrogate testing (as in e.g., in PyBioS [49]). Indeed, if the reader has a look at the associated code of Fig. 3 in appendix S1, performing timeseries surrogates significance testing by combining TimeseriesSurrogates.jl and ComplexityMeasures.jl is as seamless as if they were one package. DynamicalSystems.jl also has a component for fractal dimensions [2, Ch. 5]. Fractal dimensions themselves are complexity measures, and were initially part of ComplexityMeasures.jl. They were split off due to their codebase becoming extensive as well as being dedicated to a review article on fractal dimensions [50]. Finally, a component of DynamicalSystems.jl is about estimating optimal parameters for delay coordinate embeddings [2, Ch. 6], including the latest methods in the literature [51]. This can be used seamlessly to estimate optimal delay time and/or embedding dimension, which are crucial for the majority of complexity measures. Additionally, DelayEmbeddings.jl is used by ComplexityMeasures.jl to perform the actual delay embedding, which is a benefit since DelayEmbeddings.jl has been optimized for delay embeddings all the way down to machine code.
@@
A software based on ComplexityMeasures.jl is CausalityTools.jl [39], which implements measures for relational (cross-variable) association quantification. Relational association quantification is in itself is a huge research field, extending far beyond the scope of ComplexityMeasures.jl, where we exclusively deal with quantifying complexity within a single dataset, not between datasets. Many of these cross-variable (conditional) measures can be expressed in terms of single-variable complexity measures. For example, conditional mutual information (CMI) can be decomposed as a sum of four marginal entropy terms computed from some subset of the joint variables considered for the relational association analysis. Each entropy estimator compatible with multivariate input data is therefore also a CMI estimator. We have been very deliberate in the design of ComplexityMeasures.jl to mimic the mathematical identities between, for example, CMI and entropy. In the case of CMI, any probabilities estimator or entropy definition added to ComplexityMeasures.jl automatically enables a corresponding CMI estimator upstream in CausalityTools.jl. This enables limitless extensibility with no additional coding effort. We leave a detailed review of CausalityTools.jl and its tools for relational/association analysis for a future paper.
@@ 2.6 Performance optimizations
A large part of the development time of ComplexityMeasures.jl has been spent exclusively on optimizing the software performance. To document every single optimization we did, such as removing memory allocations, operation order re-sequencing, parallelization, etc., would take several pages and would bring us out of scope. Here we will give only a couple of characteristic examples of performance optimization via algorithmic choices. One example is the usage of Lehmer code to cast a timeseries into a sequence of ordinal patterns [52], a step required to estimate the permutation entropy. Another example is the invention of a new advanced algorithm to estimate the histogram of a multidimensional dataset whose memory requirement does not scale exponentially with the dataset dimension. The estimation is used when any sort of histogram of a delay-embedded timeseries or symbol sequence is required. The algorithm is described in Appendix A, Sec. 1 of [50].
As becomes evident in the comparison with other software (Sec. 4), these performance optimizations and algorithmic choices, along with in general following good development practices for Julia code, make ComplexityMeasures.jl by far the most performant software for complexity measure estimation, sometimes $>1,000$x faster than the competition (Table 1).
@@ 3.1 Simple complexity analysis of stock market timeseries
Here we present a straightforward analysis of stock market timeseries from the lens of complexity measures. This is an emerging topic in the literature, with relevant publications emerging only in the last 5 years. For example, Refs. [53, 54] show that regularity of the stock market anti-correlates with sample entropy. This example highlights how simple it is to integrate ComplexityMeasures.jl within a realistic data analysis workflow, as well as how many little lines of code the user needs to write: at most 1 line of code per complexity measure estimated.
For this analysis, we used a range of complexity measures: sample entropy [13], approximate entropy [12], permutation entropy [10], dispersion entropy [15], and spectral entropy [44]. These measures were estimated for the 500 stocks that compose the S&P500 index, using daily-resolution timeseries of stock closing prices for years 2000 to 2020. The complexity measures were then compared with the relative success of each stock with respect to S&P500, that is, the total price change of the stock divided by the total price change of S&P500 from start to end of the time interval (Fig.  2). The analysis shows that sample, approximate, and dispersion entropy anti-correlate with stock success, spectral entropy correlates with stock success, and permutation entropy has no relation with stock success.
This example is a modification of a group project performed by BSc Mathematics students over the course of 4 weeks. The students did not have any prior knowledge of the Julia programming language, nor of the concepts of complexity measures. This is a testament of how simple it is to learn and use ComplexityMeasures.jl, even if one has to learn an entirely new programming language from scratch.
@@ 3.3 Paintings in the complexity entropy plane
Reproducing [58]’s ordinal pattern based analysis a 2x2 square stencil (Fig. 4, left panel), supplementing with a custom non-square 4-pixel stencil using a dispersion patterns outcome space (Fig. 4, right panel), we find that artistic styles differ dramatically in their $H$-$C$ values depending on stencil pattern and outcome space, even when there are only small differences in the stencil pattern. Since [58]’s paper is not accompanied by code, we can’t actually determine whether their $H$-$C$ values are exactly reproduced here, or if any differences are caused by the addition of more paintings to the database since their analysis, differing code implementations, or other factors. However, using our software, it is trivial to further explore these topics further using different outcome space, probabilities estimators and entropy definitions/estimators, which we leave for future work.
@@ 4 Comparison with alternative software
There is a multitude (100+) of software that implement some form of complexity measures. An excellent summary of some of these software is given in the supplementary material of [60]. The overwhelming majority of these software only provide a dozen or so complexity measures as individual functions, and/or focus on a particular type of timeseries (e.g. physiological, EEG, or ECG timeseries). While we may have missed something, after a brief overview of these software, none of them appear to provide a composable orthogonal design like ComplexityMeasures.jl. Here we decided to compare more extensively with some software that have an associated peer-reviewed publication and appear to have a decent number of features: EntropyHub [61], CEPS (complexity and entropy in physiological timeseries) [6], and PyBioS [49].
In Table 1 we showcase an extensive comparison between ComplexityMeasures.jl and these three software. Explanations to superscripts in the table are as follows:
@@
-2.25in0in ComplexityMeasures.jl v3.5 EntropyHub v2 CEPS v2 PyBioS (no v) Software and development aspects Language Julia Julia, Python,
@@
Tests1 89% coverage no tests no tests no tests Extendable design Compiled documentation Introductory tutorial Explanation of measures
@@
1,530 (Appendix S2) 38 74 7
Outcome spaces 13 0 0 0 Probabilities estimators 4 1 1 1 Information definitions4 11 0 0 0 Entropy estimators 7 1 1 1 Fractal dimensions 11 0 17 0 Normalized forms Multiscale forms Performance comparison5 Permutation entropy 0.56 ms, 470.47 KiB 7.16 ms, 8.89 MiB - - Sample entropy 0.5 s, 1.71 MiB 5.23 s, 7.63 GiB - - Cosine similarity entropy 1.74 ms, 2.04 MiB 5.03 s, 4.48 GiB - - Dispersion entropy 1.46 ms, 525.75 KiB 78.55 ms, 8.23 MiB - - 30-D value histogram 1.47ms, 177.70 KiB $\infty$ (out of memory) - -
Table 1: Comparison table across software for entropic and complexity timeseries analysis. The symbols mean: ✓= has aspect, ✗= does not have aspect, ◗= partially has aspect. The numeric superscripts in the first column correspond to more extensive descriptions that we provide in the main text of Sec. 4.
@@
    Surprisingly, even though all other software have been published through peer review in reputable journals, none of them has any publicly accessible test suite that confirms the correctness of the software. ComplexityMeasures.jl has an extensive publicly accessible test suite covering 9̃0% of the total source code of the software. Without (public) tests, the only way to check for software validity is either for every user to create their own test suite, or to “just trust the developers”. Either option dramatically reduces software reliability.
@@
    This row refers to the definition of “true” or axiomatic entropies. I.e., permutation entropy does not count as a new entropy definition. ComplexityMeasures.jl is unique in this approach, as it allows the concept of entropy definition. CEPS, while it allows computing the Tsallis entropy and the Tsallis permutation entropy, it doesn’t allow composing the Tsallis entropy definition with arbitrary probability mass functions.
@@
    For the performance comparison, we evaluated the permutation, sample, cosine similarity, and dispersion entropy of a white nose timeseries with embedding dimension $m=4$, and the Shannon entropy of the histogram of a 30-dimensional chaotic Lorenz-1996 (Ref. [56]) timeseries. We report the computation time and allocated memory, showing that ComplexityMeasures.jl is routinely 10-1000x faster. For CEPS and PyBioS we cannot measure performance straightforwardly, because they are GUI-based. We note that even a difference of 10x in performance can already have a massive impact; typically one wants to estimate a complexity measure for many input timeseries (like in Sec. 3.3), or for thousands of surrogates of an input timeseries (like in Sec. 3.2).

### v3

@@ 2.3 Clarifying and educative approach to naming
The literature is full of complexity measures that are named similarly, but represent fundamentally different concepts. For example, the term “bubble entropy” [19] is not an entropy per se, but a scaled difference between two Rényi [49] entropies which have been computed by discretizing the input data in a particular manner which has to do with the bubble sort algorithm. Rényi entropy, however, is an axiomatically well-founded entropy. Similarly, permutation entropy [11] or wavelet entropy [12] are not distinct forms of entropy in terms of the mathematical definition of entropy. They are just the Shannon entropy computed by discretizing the input data into outcomes based on permutation patterns or wavelet coefficients, respectively. In fact, it is common for papers that introduce new discretization procedures, i.e. outcome spaces, to name these as some sort of “entropy” [50, 17, 16].
We believe that this abuse of terminology can be confusing, especially for newcomers to the field. Indeed, while teaching nonlinear timeseries analysis, we experience that students often interpret the Shannon entropy and permutation entropy as two different quantities. With ComplexityMeasures.jl and this paper, we aim to clarify this naming confusion, while also highlighting what the common elements between different complexity measures are: estimating probabilities from data. That is why in the software we do not promote function names like “permutation entropy”. That being said, we understand that some terms (like the permutation entropy) are very well recognized in the field and should be more accessible. Thus, for a few handpicked complexity measures we provide a shorter syntax, such as the function entropy_permutation(input; m = 3). Nevertheless we make it clear in the documentation of these “convenience functions” that they do not provide a genuinely new entropy even if named as such.
@@ 2.4 Software quality
ComplexityMeasures.jl was implemented following best practices in scientific code [44]. The software is free and open source (MIT-licensed), hosted on GitHub, and also available through the Julia package manager. The repository is composed in total of 3,841 lines of source code, 2,425 lines of test code, and 6,020 lines of documentation text according to PackageAnalyzer.jl [51]. The online documentation is very extensive [46]. It features a central tutorial; full API reference listing outcome spaces, entropy/complexity measures, and estimators provided by the software; more than a dozen of individual examples that are created by showing real code tied with its output; a developer’s documentation for contributing more features to the library. The documentation also cites all research articles introducing the implemented measures/estimators via BiBTeX, and provides an explanation and/or description of each measure/estimator implemented, making it straightforward to understand what this measure is and what it estimates without needing to consult the research article. The software is extensively tested, with coverage of 9̃0%, which means that at least 90% of the source code lines are explicitly called in the test suite. Both tests and documentation are run through continuous integration upon every committed change to the software. ComplexityMeasures.jl is continuously improved and follows agile development practices [52]. A new feature, or even the smallest bugfix, is immediately released as a new software version, which the users can obtain instantly via a standard update command provided by the Julia language. The quoted numbers in this subsection refer to version v3.7 of the software.
@@ 2.5 Part of a greater whole
ComplexityMeasures.jl can be used as a standalone software. However, it is also part of a greater whole. It is a component of the DynamicalSystems.jl [53] software library for nonlinear dynamics and nonlinear timeseries analysis, and is also the basis for the Associations.jl library for relational (or causal) timeseries analysis [45]. In this way, ComplexityMeasures.jl integrates with a wider ecosystem for timeseries and data analysis.
@@
For example, it can be immediately used with TimeseriesSurrogates.jl [54] to perform surrogate analysis testing for nonlinearities, thus eliminates the need to re-implement surrogate testing (as in e.g., in PyBioS [55]). Indeed, if the reader has a look at the associated code of Fig. 3 in Appendix S1, performing timeseries surrogates significance testing by combining TimeseriesSurrogates.jl and ComplexityMeasures.jl is as seamless as if they were one package. DynamicalSystems.jl also has a component for fractal dimensions [2, Ch. 5]. Fractal dimensions themselves are complexity measures, and were initially part of ComplexityMeasures.jl. They were split off due to their codebase becoming extensive as well as being dedicated to a review article on fractal dimensions [56] (hence, in Table. 1 we include these fractal dimensions as part of the comparison). Finally, a component of DynamicalSystems.jl is about estimating optimal parameters for delay coordinate embeddings [2, Ch. 6], including the latest methods in the literature [57]. This can be used seamlessly to estimate optimal delay time and/or embedding dimension, which are crucial for the majority of complexity measures. Additionally, DelayEmbeddings.jl is used by ComplexityMeasures.jl to perform the actual delay embedding, which is a benefit since DelayEmbeddings.jl has been optimized for delay embeddings all the way down to machine code.
@@
A software based on ComplexityMeasures.jl is Associations.jl [45], which implements measures for relational (cross-variable) association quantification. Relational association quantification is in itself is a huge research field, extending far beyond the scope of ComplexityMeasures.jl, where we exclusively deal with quantifying complexity within a single dataset, not between datasets. Many of these cross-variable (conditional) measures can be expressed in terms of single-variable complexity measures. For example, conditional mutual information (CMI) can be decomposed as a sum of four marginal entropy terms computed from some subset of the joint variables considered for the relational association analysis. Each entropy estimator compatible with multivariate input data is therefore also a CMI estimator. We have been very deliberate in the design of ComplexityMeasures.jl to mimic the mathematical identities between, for example, CMI and entropy. In the case of CMI, any probabilities estimator or entropy definition added to ComplexityMeasures.jl automatically enables a corresponding CMI estimator upstream in Associations.jl. This enables limitless extensibility with no additional coding effort. We leave a detailed review of Associations.jl and its tools for relational/association analysis for a future paper.
@@ 2.6 Performance optimizations
A large part of the development time of ComplexityMeasures.jl has been spent exclusively on optimizing the software performance. Much of the optimization stems from performant algorithmic choices that could be done in any programming language. Due to the sheer amount of algorithms included in ComplexityMeasures.jl, listing each one of this is unfeasible. Instead, we provide a few examples of such choices in the comparison section (Sec.  4).
The rest of the performance optimizations we employ are what we believe to be standard recommendations for writing performant Julia code, and most of them are discussed in the Julia manual. In summary, we have employed the following strategies.
  • Ensured type stability in all function and type definitions. Type stability allows the Julia compiler to generate highly optimized machine code, because it can determine the exact type of every variable at compile time, significantly reducing overhead and increasing performance.
  • Minimized memory allocations throughout the code base. Memory allocations can be significant performance bottlenecks in scientific computing. We have attempted to eliminate unnecessary allocations by pre-allocating storage containers (e.g. Arrays) and reusing existing memory structures where possible. As a result, memory allocations were reduced by orders of magnitude compared to competing implementations (Table. 1). Reduction in memory allocations directly translates to faster execution times and better scalability for larger datasets.
  • Separated performance-critical code into dedicated functions. This technique, known as function barriers, allows the compiler to optimize each critical section independently and make better inlining decisions.
  • Utilized StaticArrays.jl for fixed-size vector operations. Static arrays are stack-allocated rather than heap-allocated, providing immediate performance benefits through better cache utilization and zero garbage collection overhead. Additionally, their fixed size allows the compiler to perform aggressive optimizations including SIMD vectorization and loop unrolling.
  • Delegated performance-critical algorithmic steps to highly optimized and specialized Julia packages. For example, we use Distances.jl for metric space computations and DelayEmbeddings.jl for delay embeddings.
Finally, we purposely did not parallelize anything in ComplexityMeasures.jl. It is simply more efficient, and simpler from a source code perspective, to parallelize at the high level (such as when looping over different parameters or input timeseries or surrogate timeseries). Indeed, TimeseriesSurrogates.jl, a companion software discussed in §2.5, is parallelized via multithreading at a higher level.
As becomes evident in the comparison with other software (Sec. 4), these performance optimizations and algorithmic choices, along with generally following good development practices for Julia code, make ComplexityMeasures.jl by far the most performant software for complexity measure estimation, sometimes $>1,000$x faster than the competition (Table 1).
@@ 3.1 Simple complexity analysis of stock market timeseries
Here, we present a straightforward analysis of stock market timeseries from the lens of complexity measures. This is an emerging topic in the literature, with relevant publications emerging only in the last 5 years. For example, Refs. [58, 59] show that regularity of the stock market anti-correlates with sample entropy. This example highlights how simple it is to integrate ComplexityMeasures.jl within a realistic data analysis workflow, as well as how few lines of code the user needs to write: at most 1 line of code per complexity measure estimated.
For this analysis, we used a range of complexity measures: sample entropy [14], approximate entropy [13], permutation entropy [11], dispersion entropy [16], and spectral entropy [50]. These measures were estimated for the 500 stocks that compose the S&P500 index, using daily-resolution timeseries of stock closing prices for the years 2000 to 2020. The complexity measures were then compared with the relative success of each stock with respect to S&P500, that is, the total price change of the stock divided by the total price change of S&P500 from start to end of the time interval (Fig.  2). The analysis shows that sample, approximate, and dispersion entropy anti-correlate with stock success, spectral entropy correlates with stock success, and permutation entropy has no relation with stock success. The code for this analysis is also in Appendix S1 Code snippet for Figure 3..
This example is a modification of a group project performed by BSc Mathematics students over the course of 4 weeks. The students did not have any prior knowledge of the Julia programming language, nor of the concepts of complexity measures. This is a testament to how simple it is to learn and use ComplexityMeasures.jl, even if one has to learn an entirely new programming language from scratch.
@@ 3.3 Paintings in the complexity entropy plane
Reproducing [63]’s ordinal pattern based analysis a 2x2 square stencil (Fig. 4, left panel), supplementing with a custom non-square 4-pixel stencil using a dispersion patterns outcome space (Fig. 4, right panel), we find that artistic styles differ dramatically in their $H$-$C$ values depending on stencil pattern and outcome space, even when there are only small differences in the stencil pattern. Since [63]’s paper is not accompanied by code, we cannot actually determine whether their $H$-$C$ values are exactly reproduced here, or if any differences are caused by the addition of more paintings to the database since their analysis, differing code implementations, or other factors. However, using our software, it is trivial to further explore these topics further using different outcome space, probabilities estimators, and entropy definitions/estimators, which we leave for future work.
@@ 4 Comparison with alternative software
There is a multitude (100+) of software that implement some form of complexity measures. An excellent summary of some of these software is given in the supplementary material of [10]. However, the overwhelming majority of these software only provide a dozen or so complexity measures as individual functions, and/or focus on a particular type of timeseries (e.g. physiological, EEG, or ECG timeseries). While we may have missed something, after a brief overview of these software, none of them appear to provide a composable orthogonal design like ComplexityMeasures.jl.
Here we compare the performance of ComplexityMeasures.jl with notable other softwares that not only offer complexity quantification, but also have an associated peer-reviewed publication and appear to have a decent number of features: EntropyHub [65], CEPS (complexity and entropy in physiological timeseries) [6], and PyBioS [55].
In Table 1 we showcase an extensive comparison between ComplexityMeasures.jl and these three software. Explanations to the superscripts in the table are as follows:
@@
-2.25in0in ComplexityMeasures.jl v3.7 EntropyHub v2 CEPS v2 PyBioS (no v) Software and development aspects Language Julia (and Python, R, C,
due to interoperability)
Julia, Python,
@@
Tests1 92% coverage no tests no tests no tests Extendable design Compiled documentation Introductory tutorial Explanation of measures
@@
1,638 (Appendix S2) 26 74 7
Outcome spaces 13 0 0 0 Probabilities estimators 4 1 1 1 Information definitions4 12 0 0 0 Entropy estimators 7 1 1 1 Fractal dimensions 11 0 17 0 Normalized forms Multiscale functions 2 4 1 0 Performance comparison5 Permutation entropy 0.56 ms, 470.47 KiB 7.16 ms, 8.89 MiB - - Sample entropy 0.5 s, 1.71 MiB 5.23 s, 7.63 GiB - - Cosine similarity entropy 1.74 ms, 2.04 MiB 5.03 s, 4.48 GiB - - Dispersion entropy 1.46 ms, 525.75 KiB 78.55 ms, 8.23 MiB - - 30-D value histogram 1.47ms, 177.70 KiB $\infty$ (out of memory) - -
Table 1: Comparison table across software for entropic and complexity timeseries analysis. The symbols mean: ✓= has aspect, ✗= does not have aspect, ◗= partially has aspect. The numeric superscripts in the first column correspond to more extensive descriptions that we provide in the main text of Sec. 4. The codebase that produced the benchmarks can be found online in [66]. Benchmarks were run on a laptop with 11th Gen Intel(R) Core(TM) i7-1165G7 at 2.80 GHz.
@@
    Surprisingly, even though all other software have been published through peer review in reputable journals, none of them has any publicly accessible test suite that confirms the correctness of the software. ComplexityMeasures.jl has an extensive publicly accessible test suite covering $\sim$ 90% of the total source code of the software. Without (public) tests, the only way to check for software validity is either for every user to create their own test suite, or to “just trust the developers”. Either option dramatically reduces software reliability.
@@
    This row refers to the definition of “true” or axiomatic entropies. That is, permutation entropy does not count as a new entropy definition. ComplexityMeasures.jl is unique in this approach, as it allows the concept of entropy definition. CEPS, while it allows computing the Tsallis entropy and the Tsallis permutation entropy, it doesn’t allow composing the Tsallis entropy definition with arbitrary probability mass functions.
@@
    For the performance comparison, we evaluated the permutation, sample, cosine similarity, and dispersion entropy of a white noise timeseries with embedding dimension $m=4$, and the Shannon entropy of the histogram of a 30-dimensional chaotic Lorenz-1996 (Ref. [61]) timeseries. We report the computation time and the allocated memory, showing that ComplexityMeasures.jl is routinely 10-1000x faster. For CEPS and PyBioS we cannot measure performance straightforwardly, because they are GUI-based. We note that even a difference of 10x in performance can already have a massive impact in real-world applications; typically one wants to estimate a complexity measure for many input timeseries (like in Sec. 3.3), or for thousands of surrogates of an input timeseries (like in Sec. 3.2).
4.1 Comments on performance differences
We provide here some brief comments illustrating likely causes for the large performance differences between our software and the closest alternative written in the same programming language: EntropyHub. Some of the performance improvements in ComplexityMeasures.jl come from following good practices in writing performant Julia code that we discussed in Sec. 2.6. For example, for the cosine similarity entropy, compared to EntropyHub, we achieve a 2891x speedup, resulting primarily from 2196x less memory allocations. These occur largely due to the usage of the optimized Julia package DelayEmbeddings.jl for the delay embedding step. It uses optimized code generation based on statically-sized stack-allocated arrays so that the whole delay embedding operation is unrolled by the compiler. Our sample entropy implementation has 4462x less memory allocations than for EntropyHub’s implementation, resulting in a 10x faster runtime. This comes likely due to better memory management and re-using existing arrays. In addition we are using KD-trees for local neighbors searches rather than using naive brute-force distance searches.
Additional improvements in performance come from the use of different and more efficient algorithms to estimate the same quantity. For permutation entropy, we get significant speed-ups due to efficient encoding of permutation patterns into integers, which are more efficient to use for computations [^3] In general, integer encodings result in faster computations throughout the package, not only because operations on integers are simpler, but because we drastically reduce memory allocations for large input datasets compared to operating directly on arbitrary-dimensional state vectors.. This is called the “Lehmer code” [67]. Histogram-based methods are so fast in ComplexityMeasures.jl not only due to standard Julia code optimization, but also because we use a new algorithm for estimating the histogram of high-dimensional datasets which does not store bins. The memory requirement of this algorithm does not scale exponentially with the dataset dimension as is typically the case. This algorithm is used when any sort of histogram of a delay-embedded timeseries or symbol sequence is required, which is in fact the majority of methods in the library. For example, our dispersion entropy algorithm is 53x faster than EntropyHub’s implementation because it uses this histogram optimization algorithm. The algorithm is described in detail in Appendix A, Sec. 1 of [56].
These discussions illustrate how a careful approach to the software performance, in terms of following best practices, choosing the best algorithms, or even inventing new ones, can add up to overall large speed ups.